# Enhanced Duality in 4D Supergravity

### v1

@@ Abstract
U-duality imposes strong constraints on the structure of divergences in supergravity. But Gaillard-Zumino Sp$(2n_{v},\mathbb{R})$ duality in 4D, where $n_{v}$ is the number of vectors, has more symmetries than U-duality. For example, in maximal supergravity, dimension of Sp(56) is 1596, whereas its U-duality subgroup $E_{7(7)}$  has dimension 133. In comparison, D $>4$ maximal dualities are U-dualities. We argue that the extra dualities in 4D, enhancing U-duality, determine the properties of perturbative quantum supergravity. The presence/absence of enhanced dualities suggests a possible explanation of known amplitude loop computations in D-dimensional supergravities.
@@ 1 Introduction
$D=4:\qquad\rm{dim}\,[{\rm maximal\,duality}]=\rm{dim}[Sp(2n_{v},\mathbb{R})]\gg\rm{dim}[G_{U}]$ (2)
$D>4:\qquad\rm{dim}\,[{\rm maximal\,duality}]=\rm{dim}[G_{U}]$ (3)
@@
It is not widely recognized that there are two different types of supergravities in dimension D, with the same amount of local supersymmetry. Here and in WIP1 ; WIP2 we call “supergravities of type I” the well-known supergravities with local $H_{D}$ symmetry and global $E_{D-11(D-11)}$ U-duality. We call “supergravities of type II” the ones derived from higher dimensions without dualization. Their manifest local and global symmetry groups are smaller, inherited from higher dimensions.
@@ 2 UV divergences and broken local H and global G symmetries
The data has revealed a universal feature for all maximal $D>4$ supergravities: UV divergence appear at the loop order below critical. Thus, all UV divergences in maximal $D>4$ supergravities imply that the local H-symmetry and global G-symmetry are broken: the superinvariants supporting these UV divergences are given by the superspace integrals with the volume of integration which is a subspace of the total supervolume Kallosh:2023css .
The loop computations tell us that so far no UV divergences in 4D at $\mathcal{N}\geq 5$ below $L_{cr}$ were found, for example in $\mathcal{N}=5$ there is no UV divergence at $L=4<L_{cr}=5$ Bern:2014sna . There is a striking difference between the 4D and D$>$4 in dimension of maximal dualities versus U-dualities in eqs. (2) and (3) and in eq. (10) presenting UV divergences discovered at present.
Is it possible that these extra symmetries beyond U-dualities, absent in (3) and present in (2), can explain the local H-symmetry and global G-symmetry anomalies in D$>$4 associated with UV divergences, and absence of these in 4D $\mathcal{N}\geq 5$ so far? As we will argue below, the answer to this question is positive. A detailed investigation of this issue is performed in WIP1 ; WIP2 .
@@ 3 A tale of two supergravities in dimension D
The role of the gauge fields of local H-symmetry is played by the composite scalar dependent connections. The theory I has local $H_{D}$-symmetry, the theory II has a local $H_{D+1}$-symmetry, which can be gauge-fixed in various gauges. The unitary gauges in extended supergravities are studied in WIP1 , these include symmetric and Iwasawa-type gauges. Symmetric gauges tend to preserve global H symmetry and nonlinearly realized G symmetry on-shell. Iwasawa-type gauges of D-dimensional supergravities studied in WIP1 have, in general, less symmetries since they are inherited from (D+1)-dimensional supergravities compactified on a circle.
@@
Note that both supergravity I and supergravity II actions in dimension D have the maximal amount of local supersymmetries. The on-shell global symmetry $E_{5(5)}$ is present in supergravity I, but in supergravity II only $E_{4(4)}$ is manifest, as inherited from D+1.
@@ 4 Symmetric and Iwasawa-type gauges in supergravity I
A caveat here is that not every $Sp(56,\mathbb{R})$ GZ transformation leads to a different Lagrangian, there is a condition. Some of these new Lagrangians, although they might look different from the original ones, can be brought to their original form using the local change of variables, which are integration variables in the path integral. Therefore the construction of new symplectic frames in 4D in deWit:2002vt requires a GZ transformation modulo local redefinition of the fields in the Lagrangian, scalars and vectors. The remaining generators of $Sp(2n_{v},\mathbb{R})$ GZ transformation form a double quotient
$E^{{4D}}=G_{U}(\mathbb{R})\backslash Sp(2n_{v},\mathbb{R})/GL(n_{v},\mathbb{R})$ (13)
@@
$E_{{}_{\mathcal{N}=8}}^{{4D}}=E_{7(7)}(\mathbb{R})\backslash Sp(56,\mathbb{R})/GL(28,\mathbb{R})$ (14)
@@
In 6D and 8D, there are also supergravities I and supergravities II, and it is possible to choose a symmetric or an Iwasawa-type gauge in supergravity I when gauge-fixing local H-symmetry WIP1 . For example in 6D in supergravity I Tanii:1984zk there is a local $SO(5)\times SO(5)$ symmetry but only one orthogonal frame is known, the one with $GL(5,\mathbb{R})$ symmetry. The double quotient analog of the one in 4D in eq. (14) is trivial
$E_{{}_{\mathcal{N}=8}}^{6D}=E_{5(5)}(\mathbb{R})\backslash SO(5,5,\mathbb{R})/G_{v,t}(\mathbb{R})$ (16)
@@ 5 Hamiltonian path integral and quantum equivalence in 4D
$e^{-1}\mathcal{L}_{vector}=-\frac{1}{4}\mathcal{I}_{IJ}(\phi)F^{I}_{\mu\nu}F^{J\mu\nu}+\frac{1}{8}\mathcal{R}_{IJ}(\phi)\,\varepsilon^{\mu\nu\rho\sigma}F^{I}_{\mu\nu}F^{J}_{\rho\sigma}$ (17)
@@
$\left(\begin{array}[]{c}F\\ G\end{array}\right)^{\prime}=\left(\begin{array}[]{cc}A&B\\ C&D\end{array}\right)\left(\begin{array}[]{c}F\\ G\end{array}\right)$ (18)
@@
${\cal M}=\left(\begin{array}[]{cc}\mathcal{I}+\mathcal{R}\mathcal{I}^{-1}\mathcal{R}&\,\,\,-\mathcal{R}\mathcal{I}\\ -\mathcal{I}^{-1}\mathcal{R}&\,\,\,\mathcal{I}^{-1}\\ \end{array}\right)$ (21)
@@
${\mathcal{M}}^{\prime}(\phi)=E\mathcal{M}(\phi)E^{T}$ (22)
@@
$\langle{\rm out}|S|{\rm in}\rangle=\int\exp\Big{(}{i\over h}\int d^{4}x(P_{i}^{T}\dot{Q}^{i}-P^{T}{\cal M}(\phi)P))\Big{)}\prod_{x}\delta(P^{i}-P^{i}(Q))dP_{i}(x)dQ^{i}(t)$ (23)
@@
$Q_{i}^{M}=\mathcal{A}^{M}_{i}\,,\qquad P^{i}_{M}=\Omega_{MN}\mathcal{B}^{Ni}\,,\qquad\mathcal{B}^{Ni}=\varepsilon^{ijk}\partial_{j}\mathcal{A}^{N}_{k}$ (24)
@@ 6 Summary
  • Hamiltonian path integral construction for gauge theories by Faddeev-Fradkin-Tyutin-Batalin-Vilkovisky-Henneaux Faddeev:1969su , 1969
  • 4D Gaillard-Zumino duality Gaillard:1981rj , 1981, $Sp(56)\supset$ $E_{7(7)}$
@@
  • DeWit, Hamtleben, Trigiante deWit:2002vt , 2003: symplectic frames and double quotients. Duality symmetry with non-trivial double quotients, when available, serves to prove classical equivalence of symmetric and Iwasawa-type gauges in supergravities I and II and, in this way, suggests a possibility of the absence of local H-symmetry anomalies.
  • Hamiltonian formulation of duality symmetry by Henneaux, Julia, Lekeu, and Ranjbar Henneaux:2017kbx , 2017, taking into account different symplectic frames and different coset representatives, following deWit:2002vt
  • The theoretical considerations are combined in Kallosh:2023css in 2023-2024 with the amplitude computations in D-dimensional $\mathcal{N}\geq 5$ supergravities by Bern et al Bern:2023zkg , 1998-2018
  • RK et al WIP1 ; WIP2 , very recent work on the quantization of a local H-symmetry in symmetric and Iwasawa-type gauges WIP1 and construction of the quantum path integral with manifest $Sp(2n_{v},\mathbb{R})$ symmetry WIP2 . This allowed to promote the proof of classical equivalence of various versions of D-dimensional supergravities to the quantum level
Thus, classical results in GZ duality, the existence of supergravities I and II, the existence of different symplectic frames in 4D related by GZ duality were applied to the quantization of supergravity in symmetric and Iwasawa-type gauges. The proof of gauge independence (absence of H-symmetry and G-symmetry anomalies) is possible in 4D $\mathcal{N}\geq 5$ where duality is enhanced, dim [$Sp(2n_{v})]\gg$ dim [$G_{U}$] and the double quotient deWit:2002vt in (13) is non-trivial.
In $D>4$, the maximal duality is U-duality, for example in 6D $SO(5,5)\sim E_{5(5)}$. Dualities are not enhanced, and the relevant double quotients of DeWit, Hamtleben, and Trigiante are trivial. No proof is available of the on-shell gauge equivalence of symmetric and Iwasawa-type gauges and supergravities I and II. This implies a possibility of local H- global G-anomalies consistent with amplitude computations for maximal $D>4$ supergravities. The evidence for these anomalies due to UV divergences was already presented in Kallosh:2023css .
In conclusion, the existence of enhanced dualities in $N\geq 5$ supergravities in 4D provides a possible explanation of presently known cancellations of UV divergences in 4D supergravity. These include $\mathcal{N}=8$, L=3,4,5, which are usually explained by soft scalar limit due to $E_{7(7)}$  symmetry as well as the enhanced case of $N=5$, L=4, not explained by soft scalar limit due to $E_{7(7)}$  symmetry Beisert:2010jx . Thus, enhanced dualities act behind-the-scenes in amplitude computations in all 4D $N\geq 5$ cases.
@@
1. New loop computations in D$>$4 supergravities may not bring new infomation: there is a strong evidence of a H- and G-symmetries anomalies, which makes perturbative D$>$4 supergravities inconsistent.

### v2

@@ Abstract
U-duality imposes strong constraints on the structure of divergences in supergravity. But Gaillard-Zumino Sp$(2n_{v},\mathbb{R})$ duality in D=4 has more symmetries than U-duality. For example, in $\mathcal{N}=8$ the dimension of Sp(56) is 1596, whereas its U-duality subgroup $E_{7(7)}$ has a dimension of 133. In comparison, in D $>4$ maximal dualities are U-dualities. We argue that the extra dualities in 4D, enhancing U-duality, determine the properties of perturbative quantum supergravity. The presence/absence of enhanced dualities suggests a possible explanation of the results of the amplitude loop computations in D-dimensional supergravities and of the special status of D $=4$ in this respect.
@@ 1 Introduction
$D=4:\qquad\rm{dim}\,[{\rm maximal\,duality}]=\rm{dim}[Sp(2n_{v},\mathbb{R})]\gg\rm{dim}[G_{U}]\ ,$ (2)
$D>4:\qquad\rm{dim}\,[{\rm maximal\,duality}]=\rm{dim}[G_{U}]\ .$ (3)
@@
It is not widely recognized that there are two different types of supergravities in dimension D, with the same amount of local supersymmetry. Here and in WIP1 ; WIP2 , we call “supergravities of type I” the well-known supergravities with local $H_{D}$ symmetry and global $E_{D-11(D-11)}$ U-duality. We call “supergravities of type II” the ones derived from higher dimensions without dualization. Their manifest local and global symmetry groups are smaller, inherited from higher dimensions.
@@ 2 UV divergences and broken local H and global G symmetries
The data has revealed a universal feature for all maximal $D>4$ supergravities: UV divergences appear at the loop order below critical. Thus, all UV divergences in maximal $D>4$ supergravities imply that the local H-symmetry and global G-symmetry are broken: the superinvariants supporting these UV divergences are given by the superspace integrals with the volume of integration, which is a subspace of the total supervolume Kallosh:2023css .
The loop computations tell us that so far, no UV divergences in 4D at $\mathcal{N}\geq 5$ below $L_{cr}$ were found. For example, in $\mathcal{N}=5$ there is no UV divergence at $L=4<L_{cr}=5$ Bern:2014sna . There is a striking difference between the D=4 and D$>$4 in the dimension of maximal dualities versus U-dualities in eqs. (2) and (3). The same difference we see in eq. (10) presenting UV divergences discovered at present.
Is it possible that these extra symmetries beyond U-dualities, absent in (3) and present in (2), can explain the local H-symmetry and global G-symmetry anomalies in D$>$4 associated with UV divergences, and the absence of these in 4D $\mathcal{N}\geq 5$ so far? As we will argue below, the answer to this question is positive. A detailed investigation of this issue is performed in WIP1 ; WIP2 .
@@ 3 A tale of two supergravities in dimension D
The role of the gauge fields of local H-symmetry is played by the composite scalar-dependent connections. Theory I has local $H_{D}$-symmetry, and Theory II has a local $H_{D+1}$-symmetry, which can be gauge-fixed in various gauges. The unitary gauges in extended supergravities are studied in WIP1 , these include symmetric and Iwasawa-type gauges. Symmetric gauges tend to preserve global H symmetry and nonlinearly realized G symmetry on-shell. Iwasawa-type gauges of D-dimensional supergravities studied in WIP1 have, in general, less symmetries since they are inherited from (D+1)-dimensional supergravities compactified on a circle.
@@
Note that both supergravity I and supergravity II actions in dimension D have the maximal amount of local supersymmetries. The on-shell global symmetry $E_{5(5)}$ is present in supergravity I, but in supergravity II, only $E_{4(4)}$ is manifest, as inherited from D+1.
@@ 4 Symmetric and Iwasawa-type gauges in supergravity I
A caveat here is that not every $Sp(56,\mathbb{R})$ GZ transformation leads to a different Lagrangian, there is a condition. Some of these new Lagrangians, although they might look different from the original ones, can be brought to their original form using the local change of variables, which are integration variables in the path integral. Therefore, the construction of new symplectic frames in 4D in deWit:2002vt requires a GZ transformation modulo local redefinition of the fields in the Lagrangian, scalars, and vectors. The remaining generators of $Sp(2n_{v},\mathbb{R})$ GZ transformation form a double quotient
$E^{{4D}}=G_{U}(\mathbb{R})\backslash Sp(2n_{v},\mathbb{R})/GL(n_{v},\mathbb{R})\ .$ (13)
@@
$E_{{}_{\mathcal{N}=8}}^{{4D}}=E_{7(7)}(\mathbb{R})\backslash Sp(56,\mathbb{R})/GL(28,\mathbb{R})\ .$ (14)
@@
In 6D and 8D, there are also supergravities I and supergravities II, and it is possible to choose a symmetric or an Iwasawa-type gauge in supergravity I when gauge-fixing local H-symmetry WIP1 . For example, in 6D in supergravity I Tanii:1984zk there is a local $SO(5)\times SO(5)$ symmetry, but only one orthogonal frame is known, the one with $GL(5,\mathbb{R})$ symmetry. The double quotient analog of the one in 4D in eq. (14) is trivial
$E_{{}_{\mathcal{N}=8}}^{6D}=E_{5(5)}(\mathbb{R})\backslash SO(5,5,\mathbb{R})/G_{v,t}(\mathbb{R})\ ,$ (16)
@@ 5 Hamiltonian path integral and quantum equivalence in 4D
$e^{-1}\mathcal{L}_{vector}=-\frac{1}{4}\mathcal{I}_{IJ}(\phi)F^{I}_{\mu\nu}F^{J\mu\nu}+\frac{1}{8}\mathcal{R}_{IJ}(\phi)\,\varepsilon^{\mu\nu\rho\sigma}F^{I}_{\mu\nu}F^{J}_{\rho\sigma}\ .$ (17)
@@
$\left(\begin{array}[]{c}F\\ G\end{array}\right)^{\prime}=\left(\begin{array}[]{cc}A&B\\ C&D\end{array}\right)\left(\begin{array}[]{c}F\\ G\end{array}\right)\ ,$ (18)
@@
${\cal M}=\left(\begin{array}[]{cc}\mathcal{I}+\mathcal{R}\mathcal{I}^{-1}\mathcal{R}&\,\,\,-\mathcal{R}\mathcal{I}^{-1}\\ -\mathcal{I}^{-1}\mathcal{R}&\,\,\,\mathcal{I}^{-1}\\ \end{array}\right)\ .$ (21)
@@
${\mathcal{M}}^{\prime}(\phi)=E\mathcal{M}(\phi)E^{T}\ .$ (22)
@@
$\langle{\rm out}|S|{\rm in}\rangle=\int\exp\Big{(}{i\over h}\int d^{4}x(P_{i}^{T}\dot{Q}^{i}-P^{T}{\cal M}(\phi)P))\Big{)}\prod_{x}\delta(P^{i}-P^{i}(Q))dP_{i}(x)dQ^{i}(t)\ .$ (23)
@@
$Q_{i}^{M}=\mathcal{A}^{M}_{i}\,,\qquad P^{i}_{M}=\Omega_{MN}\mathcal{B}^{Ni}\,,\qquad\mathcal{B}^{Ni}=\varepsilon^{ijk}\partial_{j}\mathcal{A}^{N}_{k}\ .$ (24)
@@ 6 Summary
  • Hamiltonian path integral construction for gauge theories by Faddeev-Fradkin-Tyutin-Batalin-Vilkovisky-Henneaux Faddeev:1969su , 1969.
  • 4D Gaillard-Zumino duality Gaillard:1981rj , 1981, $Sp(56)\supset$ $E_{7(7)}$.
@@
  • DeWit, Samtleben, Trigiante deWit:2002vt , 2003: symplectic frames and double quotients. Duality symmetry with non-trivial double quotients, when available, serves to prove classical equivalence of symmetric and Iwasawa-type gauges in supergravities I and II and, in this way, suggests a possibility of the absence of local H-symmetry anomalies.
  • Hamiltonian formulation of duality symmetry by Henneaux, Julia, Lekeu, and Ranjbar Henneaux:2017kbx , 2017, taking into account different symplectic frames and different coset representatives, following deWit:2002vt .
  • The theoretical considerations are combined in Kallosh:2023css in 2023-2024 with the amplitude computations in D-dimensional $\mathcal{N}\geq 5$ supergravities by Bern et al Bern:2023zkg , 1998-2018.
  • RK et al WIP1 ; WIP2 , very recent work on the quantization of a local H-symmetry in symmetric and Iwasawa-type gauges WIP1 and construction of the quantum path integral with manifest $Sp(2n_{v},\mathbb{R})$ symmetry WIP2 . This allowed the promotion of the proof of classical equivalence of various versions of D-dimensional supergravities to the quantum level.
Thus, classical results in GZ duality, the existence of supergravities I and II, and the existence of different symplectic frames in 4D related by GZ duality were applied to the quantization of supergravity in symmetric and Iwasawa-type gauges. The proof of gauge independence (absence of H-symmetry and G-symmetry anomalies) is possible in 4D $\mathcal{N}\geq 5$ where duality is enhanced, dim [$Sp(2n_{v})]\gg$ dim [$G_{U}$] and the double quotient deWit:2002vt in (13) is non-trivial.
In $D>4$, the maximal duality is U-duality, for example in 6D $SO(5,5)\sim E_{5(5)}$. Dualities are not enhanced, and the relevant double quotients of DeWit, Samtleben, and Trigiante are trivial. No proof is available of the on-shell gauge equivalence of symmetric and Iwasawa-type gauges and supergravities I and II. This implies a possibility of local H- global G-anomalies consistent with amplitude computations for maximal $D>4$ supergravities. The evidence for these anomalies due to UV divergences was already presented in Kallosh:2023css .
In conclusion, the existence of enhanced dualities in $N\geq 5$ supergravities in 4D provides a possible explanation of presently known cancellations of UV divergences in 4D supergravity. These include $\mathcal{N}=8$, L=3,4,5, which are usually explained by soft scalar limit due to $E_{7(7)}$  symmetry as well as the enhanced case of $N=5$, L=4, not explained by soft scalar limit due to $E_{7(7)}$  symmetry Beisert:2010jx . Thus, enhanced dualities act behind the scenes in amplitude computations in all 4D $N\geq 5$ cases.
@@
1. New loop computations in D$>$4 supergravities may not bring new information: there is strong evidence of a H- and G-symmetries anomalies, which makes perturbative D$>$4 supergravities inconsistent.